OpenStudy (anonymous):

Write the equation in slope intercept form that contains the point (4,7) and (0,5)

OpenStudy (aroub):

first find the slope 5-7/0-4 -2/-4 1/2 y-y1=m(x-x1) , m= 1/2 , (0,5) y-5=1/2(x-0)

OpenStudy (anonymous):

aroub isn't that point slope form?

OpenStudy (aroub):

yes!! :)

OpenStudy (aroub):

oh lol!! totally forgot!! soorrrryyyyy!!!! :(

OpenStudy (aroub):

so the answer to ur question is y=1/2x +5 right??

OpenStudy (anonymous):

slope intercept form would be y=.5x+5

OpenStudy (aroub):

yup same thing!! :D

OpenStudy (anonymous):

yup!